# Time-dependent Regularized 13-Moment Equations with Onsager Boundary Conditions in the Linear Regime

### v1

@@ 1. Introduction
Instead of directly solving the Boltzmann equation, many researchers focus on the macroscopic moments of the distribution function and modify classical fluid equations, such as the Navier-Stokes equations, to describe the mechanics of moderately rarefied gases. Two prominent techniques used to derive macroscopic fluid equations from the Boltzmann equation are the Grad’s moment method in [14] and the Chapman-Enskog expansion in [10]. Despite their successes in capturing near-equilibrium effects, these methods suffer from instability and the loss of hyperbolicity, particularly when higher-order expansions are adopted or more moments are included. Recent researches address these limitations by employing various regularization techniques ([18, 21, 30, 4, 23, 15]) to enhance the accuracy and stability of the methods across a broader range of rarefied gas dynamics.
In this work, we will investigate the regularized version of Grad’s 13-moment equations proposed in [30], which can be interpreted as a Chapman–Enskog like procedure applied to the Grad’s 13 moments. This regularization, abbreviated as R13, has been developed for both Maxwell ([28]) and non-Maxwell molecules ([31, 9]) and validated across diverse applications ([32, 25, 33]). The R13 equations are attractive since the equations include only second-order derivatives while attaining the super-Burnett order. Here the super-Burnett order means that the super-Burnett equations, which are the result of the Chapman-Enskog expansion up to the third order, can be derived from the R13 equations. Note that the original super-Burnett equations contain fourth-order derivatives, causing significant difficulties in the numerical discretization and formulation of boundary conditions. Also, the R13 equations are equipped with reliable boundary conditions ([35, 29]) and have been shown to be robust in numerical simulations ([24]). Recently in [34], Onsager boundary conditions are proposed to ensure the stability of boundary value problems of linear R13 equations for Maxwell molecules. However, it is not straightforward to generalize these Onsager boundary conditions to non-Maxwell molecules, even in the linear case. The main reason is the loss of a symmetric structure required by the Onsager boundary conditions (see [5]). To remedy this, in [8], the linear R13 equations for arbitrary elastic collision models were re-derived using a different method. The new derivation relies on the simpler form of the Chapman-Enskog expansion for steady-state equations compared with the time-dependent case, so that it cannot be applied directly to the time-dependent case.
@@
where $\mathcal{L}$ is a self-adjoint negative semidefinite linear operator; $x_{k}$ and $\xi_{k}$ are position and velocity variables, respectively; $\mathrm{Kn}$ is the Knudsen number defined as the ratio of the mean free path of the molecules to macroscopic length scale. Note that here we have used the summation convention: when the same index appears twice, it is assumed that the sum of applied for this index, and the range is from 1 to 3. In order that the Onsager boundary conditions can be applied, the linear moment equations should take the form
@@
where $\mathbf{A}_{0}$ is the symmetric mass matrix, $\mathbf{A}_{k}$ is the symmetric discretization of operator $\xi_{k}$, and $\mathbf{L}$ is the discretization of operator $\mathcal{L}$ which is symmetric negative semidefinite. The matrix $\mathbf{A}_{0}$ is allowed to have zero eigenvalues to cover the parabolic scenarios such as the R13 equations.
The $L^{2}$ stability of (2) in a bounded domain $\Omega$ requires appropriate boundary conditions. This can be seen by multiplying $\boldsymbol{u}^{\intercal}$ on (2) and integrating over $\Omega$, yielding
@@
where $\boldsymbol{n}=(n_{1},n_{2},n_{3})^{\intercal}$ denotes the outward normal vector at boundary $\partial\Omega$. Hence, the $L^{2}$ stability can be achieved if the flux across the boundary is bounded, i.e.,
@@
for a given matrix $\mathbf{M}$ and an external source vector $\boldsymbol{g}_{\mathrm{ext}}=\boldsymbol{g}_{\mathrm{ext}}(t,\boldsymbol{x})$ from boundary conditions. With the above stability, the uniqueness of the solution of (2) can be achieved.
@@ 2.1. Choice of variables
The 13-moment equations were first proposed by [14], where the distribution function was expanded into an infinite series using Hermite polynomials. The coefficients of the polynomials were regarded as moments, and Grad chose to include the first 13 moments, namely the density ($\rho$), momentum ($\rho v_{i}$), pressure ($\rho\theta$), stress tensor ($\sigma_{ij}$) and heat flux ($q_{i}$), into his equations. In later developments of 13-moment systems, such a choice of the 13 moments becomes a standard and is followed by most works [30, 22, 36]. In particular, for the regularized 13-moment equations for Maxwell molecules, such a choice of moments matches exactly with the analysis using the order of magnitude method [28, Chapter 8]. It is seen that for linear equations, these moments, denoted by $\langle\boldsymbol{\phi}_{13},f\rangle$ with $\boldsymbol{\phi}_{13}$ being a 13-dimensional vector of polynomials of $\boldsymbol{\xi}$, fully describes the distribution function $f$ up to order $O(\mathrm{Kn})$, which means for any polynomial $r(\boldsymbol{\xi})$ satisfying $\langle\boldsymbol{\phi}_{13},r\rangle=0$, the moment $\langle r,f\rangle$ has at least order $O(\mathrm{Kn}^{2})$ when performing asymptotic analysis. This property leads to the super-Burnett order of R13 equations for Maxwell molecules, meaning that the super-Burnett equations can be derived from the R13 equations by Chapman-Enskog expansion. However, it is also known that this no longer holds for other types of molecules [27]. Inspired by this fact, in our study of 13-moment equations for general gas molecules, we will adopt a different set of variables $\langle\tilde{\boldsymbol{\phi}}_{13},f\rangle$, such that we can again obtain $\langle r,f\rangle\sim o(\mathrm{Kn})$ once $\langle\tilde{\boldsymbol{\phi}}_{13},r\rangle=0$. This will help achieve the super-Burnett order for the moment equations.
@@ 2.2. Moment equations and boundary conditions
(8) $\frac{\partial\theta}{\partial t}+\frac{2}{3}\frac{\partial v_{j}}{\partial x_{j}}+\frac{2}{3}k_{0}\frac{\partial\bar{q}_{j}}{\partial x_{j}}-k_{1}\mathrm{Kn}\frac{\partial^{2}\theta}{\partial x_{j}\partial x_{j}}+k_{2}\mathrm{Kn}\frac{\partial^{2}\bar{\sigma}_{ij}}{\partial x_{i}\partial x_{j}}=0,$
(9) $\frac{\partial v_{i}}{\partial t}+\frac{\partial\rho}{\partial x_{i}}+\frac{\partial\theta}{\partial x_{i}}-k_{3}\mathrm{Kn}\frac{\partial^{2}v_{\langle i}}{\partial x_{j\rangle}\partial x_{j}}-k_{4}\mathrm{Kn}\frac{\partial^{2}\bar{q}_{\langle i}}{\partial x_{j\rangle}\partial x_{j}}+k_{5}\frac{\partial\bar{\sigma}_{ij}}{\partial x_{j}}=0,$
$\frac{\partial\bar{q}_{i}}{\partial t}+\frac{5}{2}k_{0}\frac{\partial\theta}{\partial x_{i}}$
(10) $\quad{}-\frac{5}{2}k_{4}\mathrm{Kn}\frac{\partial^{2}v_{\langle i}}{\partial x_{j\rangle}\partial x_{j}}-2k_{6}\mathrm{Kn}\frac{\partial^{2}\bar{q}_{j}}{\partial x_{j}\partial x_{i}}-\frac{12}{5}k_{7}\mathrm{Kn}\frac{\partial^{2}\bar{q}_{\langle i}}{\partial x_{j\rangle}\partial x_{j}}+k_{8}\frac{\partial\bar{\sigma}_{ij}}{\partial x_{j}}=-\frac{1}{\mathrm{Kn}}\frac{2}{3}l_{1}\bar{q}_{i},$
$\frac{\partial\bar{\sigma}_{ij}}{\partial t}+3k_{2}\mathrm{Kn}\frac{\partial^{2}\theta}{\partial x_{\langle i}\partial x_{j\rangle}}+2k_{5}\frac{\partial v_{\langle i}}{\partial x_{j\rangle}}+\frac{4}{5}k_{8}\frac{\partial\bar{q}_{\langle i}}{\partial x_{j\rangle}}$
(11) $\quad{}-2k_{9}\mathrm{Kn}\frac{\partial^{2}\bar{\sigma}_{\langle ij}}{\partial x_{k\rangle}\partial x_{k}}-k_{10}\mathrm{Kn}\frac{\partial^{2}\bar{\sigma}_{k\langle i}}{\partial x_{j\rangle}\partial x_{k}}=-\frac{1}{\mathrm{Kn}}l_{2}\bar{\sigma}_{ij}.$
The coefficients $k_{i}$ and $l_{i}$ have been computed for gas molecules whose intermolecular potential satisfies inverse power laws. Values for these coefficients for some power indices are given in the Appendix A.
@@
$\frac{\partial\theta}{\partial t}+\frac{2}{3}\frac{\partial v_{j}}{\partial x_{j}}+\frac{2}{3}\frac{\partial q_{j}}{\partial x_{j}}=0,\qquad\frac{\partial v_{i}}{\partial t}+\frac{\partial\rho}{\partial x_{i}}+\frac{\partial\theta}{\partial x_{i}}+\frac{\partial\sigma_{ij}}{\partial x_{j}}=0,$
one can observe the relationship between $\bar{\sigma}_{ij}$, $\bar{q}_{i}$ and $\sigma_{ij}$, $q_{i}$:
$\sigma_{ij}=k_{5}\bar{\sigma}_{ij}-k_{4}\mathrm{Kn}\frac{\partial\bar{q}_{\langle i}}{\partial x_{j\rangle}}-k_{3}\mathrm{Kn}\frac{\partial v_{\langle i}}{\partial x_{j\rangle}},\quad q_{i}=k_{0}\bar{q_{i}}-\frac{3}{2}k_{1}\mathrm{Kn}\frac{\partial\theta}{\partial x_{i}}+\frac{3}{2}k_{2}\mathrm{Kn}\frac{\partial\bar{\sigma}_{ij}}{\partial x_{j}}.$
Such a relationship allows us to calculate the stress tensor and heat flux once the equations (7)–(11) are solved.
@@
$m_{46}\bar{\sigma}_{\tau_{i}n}-m_{47}\mathrm{Kn}\frac{\partial v_{\langle\tau_{i}}}{\partial x_{n\rangle}}-m_{48}\mathrm{Kn}\frac{\partial\bar{q}_{\langle\tau_{i}}}{\partial x_{n\rangle}}$
(16) $\quad=\tilde{\chi}\left[-m_{41}(v_{\tau_{i}}-v^{W}_{\tau_{i}})+m_{42}\bar{q}_{\tau_{i}}+m_{43}\mathrm{Kn}\frac{\partial\bar{\sigma}_{\tau_{i}j}}{\partial x_{j}}+m_{44}\mathrm{Kn}\frac{\partial\bar{\sigma}_{\langle\tau_{i}n}}{\partial x_{n\rangle}}-m_{45}\mathrm{Kn}\frac{\partial\theta}{\partial x_{\tau_{i}}}\right],i=1,2,$
$m_{56}\bar{\sigma}_{\tau_{i}n}-m_{57}\mathrm{Kn}\frac{\partial v_{\langle\tau_{i}}}{\partial x_{n\rangle}}-m_{58}\mathrm{Kn}\frac{\partial\bar{q}_{\langle\tau_{i}}}{\partial x_{n\rangle}}$
(17) $\quad=\tilde{\chi}\left[-m_{51}(v_{\tau_{i}}-v^{W}_{\tau_{i}})+m_{52}\bar{q}_{\tau_{i}}+m_{53}\mathrm{Kn}\frac{\partial\bar{\sigma}_{\tau_{i}j}}{\partial x_{j}}+m_{54}\mathrm{Kn}\frac{\partial\bar{\sigma}_{\langle\tau_{i}n}}{\partial x_{n\rangle}}-m_{55}\mathrm{Kn}\frac{\partial\theta}{\partial x_{\tau_{i}}}\right],i=1,2,$
$-m_{66}\bar{q}_{n}-\mathrm{Kn}\left(m_{67}\frac{\partial\bar{\sigma}_{\langle nn}}{\partial x_{n\rangle}}+m_{68}\frac{\partial\bar{\sigma}_{nj}}{\partial x_{j}}+m_{69}\frac{\partial\theta}{\partial x_{n}}\right)$
(18) $\quad=\tilde{\chi}\left[-m_{61}(\theta-\theta^{W})+m_{62}\bar{\sigma}_{nn}+m_{63}\mathrm{Kn}\frac{\partial\bar{q}_{j}}{\partial x_{j}}+m_{64}\mathrm{Kn}\frac{\partial\bar{q}_{\langle n}}{\partial x_{n\rangle}}+m_{65}\mathrm{Kn}\frac{\partial v_{\langle n}}{\partial x_{n\rangle}}\right],$
@@
In these equations, $m_{jk}$ are constants depending on the molecular interactions. As examples, we will consider inverse-power-law models in which the repulsive force between two gas molecules is proportional to the $\eta$th power of the distance between them. For such models, the values of the constants $k_{i}$ and $m_{ij}$ are tabulated in the Appendix A for some power indices.

### v2

@@ 1. Introduction
Instead of directly solving the Boltzmann equation, many researchers focus on the macroscopic moments of the distribution function and modify classical fluid equations, such as the Navier-Stokes equations, to describe the mechanics of moderately rarefied gases. Two prominent techniques used to derive higher-order (in terms of Knudsen number) macroscopic fluid equations from the Boltzmann equation are the Grad’s moment method in [17] and the Chapman-Enskog expansion in [12]. Despite their successes in capturing near-equilibrium effects, the Grad’s 13-moment equations may experience a loss of hyperbolicity and unphysical subshocks. Moreover, the second-order Burnett and third-order super-Burnett equations, which are derived from the Chapman-Enskog expansion, suffer from instability. Recent researches address these limitations by employing various regularization techniques ([21, 24, 34, 6, 26, 18]) to enhance the accuracy and stability of the methods across a broader range of rarefied gas dynamics.
In this work, we will investigate the regularized version of Grad’s 13-moment equations proposed in [34], which can be interpreted as a Chapman–Enskog like procedure applied to the Grad’s 13 moments. This regularization, abbreviated as R13, has been developed for both Maxwell ([32]) and non-Maxwell molecules ([36, 11]) and validated across diverse applications ([37, 28, 39]). The R13 equations are attractive since the equations include only second-order derivatives while attaining the super-Burnett order. Here the super-Burnett order means that the super-Burnett equations, which are the result of the Chapman-Enskog expansion up to the third order, can be derived from the R13 equations. Note that the original super-Burnett equations contain fourth-order derivatives, causing significant difficulties in the numerical discretization and formulation of boundary conditions. Due to the contribution of higher-order terms, the R13 equations can describe various rarefaction effects such as Knudsen boundary layers and heat fluxes from cold areas to hot areas ([27]). Also, the R13 equations are equipped with reliable boundary conditions ([41, 33]) and have been shown to be robust in numerical simulations ([27]). Recently in [40], Onsager boundary conditions are proposed to ensure the stability of boundary value problems of linear R13 equations for Maxwell molecules. However, it is not straightforward to generalize these Onsager boundary conditions to non-Maxwell molecules, even in the linear case. The main reason is the loss of a symmetric structure required by the Onsager boundary conditions (see [7]). To remedy this, in [10], the linear R13 equations for arbitrary elastic collision models were re-derived using a different method. The new derivation relies on the simpler form of the Chapman-Enskog expansion for steady-state equations compared with the time-dependent case, so that it cannot be applied directly to dynamic problems.
@@
where $\mathcal{L}$ is a self-adjoint negative semidefinite linear operator; $x_{k}$ and $\xi_{k}$ are position and velocity variables, respectively; $\mathrm{Kn}$ is the Knudsen number defined as the ratio of the mean free path of the molecules to macroscopic length scale. Note that here we have used the summation convention: when the same index appears twice, it is assumed that the sum of applied for this index, and the range is from 1 to 3. Due to the linearization, the model is mainly applicable for fluids with low speed such as microflows, and numerical experiments show good agreement with results obtained from the DSMC method (see [42]). In order that the Onsager boundary conditions can be applied, the linear moment equations should take the form
@@
where $\mathbf{A}_{0}$ is the symmetric mass matrix, $\mathbf{A}_{k}$ is the symmetric discretization of operator $\xi_{k}$, and $\mathbf{L}$ is the discretization of operator $\mathcal{L}$ which is symmetric negative semidefinite. The matrix $\mathbf{A}_{0}$ is allowed to have zero eigenvalues to cover the parabolic scenarios such as the linearized R13 equations, which has been used to study phenomena including evaporation [13] and rarefied effects in flows around a sphere [2]. It should be noted that the symmetric structure in the above matrices plays a key role in obtaining the second law of thermodynamics for our moment equations. The similar idea of maintaining symmetry to comply with thermodynamics has also been applied to the derivation of other moment equations ([30, 43]).
When considering thermodynamics of (2) within a bounded domain $\Omega$, the entropy production from its boundary should be taken into account. This can be seen by multiplying $\boldsymbol{u}^{\intercal}$ on (2) and integrating over $\Omega$, yielding
@@
where $\boldsymbol{n}=(n_{1},n_{2},n_{3})^{\intercal}$ denotes the outward normal vector at boundary $\partial\Omega$. Hence, the second law of thermodynamics can be achieved if the flux across the boundary is bounded, i.e.,
@@
for a given matrix $\mathbf{M}$ and an external source vector $\boldsymbol{g}_{\mathrm{ext}}=\boldsymbol{g}_{\mathrm{ext}}(t,\boldsymbol{x})$ from boundary conditions. With the above entropic stability, the uniqueness of the solution of (2) can be achieved.
@@ 2.1. Choice of variables
The 13-moment equations were first proposed by [17], where the distribution function was expanded into an infinite series using Hermite polynomials. The coefficients of the polynomials were regarded as moments, and Grad chose to include the first 13 moments, namely the density ($\rho$), momentum ($\rho v_{i}$), pressure ($\rho\theta$), stress tensor ($\sigma_{ij}$) and heat flux ($q_{i}$), into his equations. In later developments of 13-moment systems, such a choice of the 13 moments becomes a standard and is followed by most works [34, 25, 44]. In particular, for the regularized 13-moment equations for Maxwell molecules, such a choice of moments matches exactly with the analysis using the order of magnitude method [32, Chapter 8]. It is seen that for linear equations, these moments, denoted by $\langle\boldsymbol{\phi}_{13},f\rangle$ with $\boldsymbol{\phi}_{13}$ being a 13-dimensional vector of polynomials of $\boldsymbol{\xi}$, fully describes the distribution function $f$ up to order $O(\mathrm{Kn})$, which means for any polynomial $r(\boldsymbol{\xi})$ satisfying $\langle\boldsymbol{\phi}_{13},r\rangle=0$, the moment $\langle r,f\rangle$ has at least order $O(\mathrm{Kn}^{2})$ when performing asymptotic analysis. This property leads to the super-Burnett order of R13 equations for Maxwell molecules, meaning that the super-Burnett equations can be derived from the R13 equations by Chapman-Enskog expansion. However, it is also known that this no longer holds for other types of molecules [31]. Inspired by this fact, in our study of 13-moment equations for general gas molecules, we will adopt a different set of variables $\langle\tilde{\boldsymbol{\phi}}_{13},f\rangle$, such that we can again obtain $\langle r,f\rangle\sim o(\mathrm{Kn})$ once $\langle\tilde{\boldsymbol{\phi}}_{13},r\rangle=0$. This will help achieve the super-Burnett order for the moment equations.
@@ 2.2. Moment equations and boundary conditions
(7) $\frac{\partial\rho}{\partial t}+\nabla\cdot\boldsymbol{v}=0,$
(8) $\frac{\partial\theta}{\partial t}+\frac{2}{3}\nabla\cdot\boldsymbol{v}+\frac{2}{3}k_{0}\nabla\cdot\boldsymbol{\bar{q}}-k_{1}\mathrm{Kn}\Delta\theta+k_{2}\mathrm{Kn}\nabla\cdot(\nabla\cdot\boldsymbol{\bar{\sigma}})=0,$
(9) $\frac{\partial\boldsymbol{v}}{\partial t}+\nabla\rho+\nabla\theta-k_{3}\mathrm{Kn}\nabla\cdot(\nabla\boldsymbol{v})_{\mathrm{stf}}-k_{4}\mathrm{Kn}\nabla\cdot(\nabla\boldsymbol{\bar{q}})_{\mathrm{stf}}+k_{5}\nabla\cdot\boldsymbol{\bar{\sigma}}=0,$
$\frac{\partial\boldsymbol{\bar{q}}}{\partial t}+\frac{5}{2}k_{0}\nabla\theta-\frac{5}{2}k_{4}\mathrm{Kn}\nabla\cdot(\nabla\boldsymbol{v})_{\mathrm{stf}}-2k_{6}\mathrm{Kn}\nabla(\nabla\cdot\boldsymbol{\bar{q}})$
(10) $\quad{}-\frac{12}{5}k_{7}\mathrm{Kn}\nabla\cdot(\nabla\boldsymbol{\bar{q}})_{\mathrm{stf}}+k_{8}\nabla\cdot\boldsymbol{\bar{\sigma}}=-\frac{1}{\mathrm{Kn}}\frac{2}{3}l_{1}\boldsymbol{\bar{q}},$
$\frac{\partial\boldsymbol{\bar{\sigma}}}{\partial t}+3k_{2}\mathrm{Kn}\left(\nabla^{2}\theta\right)_{\mathrm{stf}}+2k_{5}(\nabla\boldsymbol{v})_{\mathrm{stf}}+\frac{4}{5}k_{8}(\nabla\boldsymbol{\bar{q}})_{\mathrm{stf}}$
(11) $\quad{}-2k_{9}\mathrm{Kn}\nabla\cdot(\nabla\boldsymbol{\bar{\sigma}})_{\mathrm{stf}}-k_{10}\mathrm{Kn}\left(\nabla(\nabla\cdot\boldsymbol{\bar{\sigma}})\right)_{\mathrm{stf}}=-\frac{1}{\mathrm{Kn}}l_{2}\boldsymbol{\bar{\sigma}}.$
Here $(\cdot)_{\mathrm{stf}}$ refers to the symmetric and trace-free part of a given tensor, which is defined entrywisely as $(\cdot)_{ij}\mapsto((\cdot)_{\mathrm{stf}})_{ij}=(\cdot)_{\langle ij\rangle}$ for a matrix and $(\cdot)_{ijk}\mapsto((\cdot)_{\mathrm{stf}})_{ijk}=(\cdot)_{\langle ijk\rangle}$ for a 3-tensor. The detailed definition can be found in [32]. The coefficients $k_{i}$ and $l_{i}$ have been computed for gas molecules whose intermolecular potential satisfies inverse power laws. Values for these coefficients for some power indices are given in the Appendix B. Although these coefficients look arbitrary, our derivation in section 3 will reveal that they are fully determined by the linearized and nondimensionalized Boltzmann collision operator, which depends only on the potential energy between two gas molecules. In inverse-power-law models, only two parameters exist: the intensity coefficient of the potential and the power $\eta$. After nondimensionalization, the intensity coefficient is integrated into the Knudsen number, and thus $\eta$ is the only parameter that dictates all the coefficients.
@@
$\frac{\partial\theta}{\partial t}+\frac{2}{3}\nabla\cdot\boldsymbol{v}+\frac{2}{3}\nabla\cdot\boldsymbol{q}=0,\qquad\frac{\partial\boldsymbol{v}}{\partial t}+\nabla\rho+\nabla\theta+\nabla\cdot\boldsymbol{\sigma}=0,$
one can observe the relationship between $\boldsymbol{\bar{\sigma}}$, $\boldsymbol{\bar{q}}$ and $\boldsymbol{\sigma}$, $\boldsymbol{q}$:
(12) $\boldsymbol{\sigma}=k_{5}\boldsymbol{\bar{\sigma}}-k_{4}\mathrm{Kn}(\nabla\boldsymbol{\bar{q}})_{\mathrm{stf}}-k_{3}\mathrm{Kn}(\nabla\boldsymbol{v})_{\mathrm{stf}},\quad\boldsymbol{q}=k_{0}\boldsymbol{\bar{q}}-\frac{3}{2}k_{1}\mathrm{Kn}\nabla\theta+\frac{3}{2}k_{2}\mathrm{Kn}\nabla\cdot\boldsymbol{\bar{\sigma}}.$
Such a relationship allows us to calculate the stress tensor and heat flux once the equations (7)–(11) are solved. We remark that the moment equations (7)–(11) we derived is Galilean invariant. The detailed proof can be found in Appendix A.
@@
$m_{46}\bar{\sigma}_{\tau_{i}n}+m_{47}\mathrm{Kn}\frac{\partial v_{\langle\tau_{i}}}{\partial x_{n\rangle}}+m_{48}\mathrm{Kn}\frac{\partial\bar{q}_{\langle\tau_{i}}}{\partial x_{n\rangle}}$
(17) $\quad=-\tilde{\chi}\left[-m_{41}(v_{\tau_{i}}-v^{W}_{\tau_{i}})+m_{42}\bar{q}_{\tau_{i}}+m_{43}\mathrm{Kn}\frac{\partial\bar{\sigma}_{\tau_{i}j}}{\partial x_{j}}+m_{44}\mathrm{Kn}\frac{\partial\bar{\sigma}_{\langle\tau_{i}n}}{\partial x_{n\rangle}}-m_{45}\mathrm{Kn}\frac{\partial\theta}{\partial x_{\tau_{i}}}\right],i=1,2,$
$m_{56}\bar{\sigma}_{\tau_{i}n}+m_{57}\mathrm{Kn}\frac{\partial v_{\langle\tau_{i}}}{\partial x_{n\rangle}}+m_{58}\mathrm{Kn}\frac{\partial\bar{q}_{\langle\tau_{i}}}{\partial x_{n\rangle}}$
(18) $\quad=-\tilde{\chi}\left[-m_{51}(v_{\tau_{i}}-v^{W}_{\tau_{i}})+m_{52}\bar{q}_{\tau_{i}}+m_{53}\mathrm{Kn}\frac{\partial\bar{\sigma}_{\tau_{i}j}}{\partial x_{j}}+m_{54}\mathrm{Kn}\frac{\partial\bar{\sigma}_{\langle\tau_{i}n}}{\partial x_{n\rangle}}-m_{55}\mathrm{Kn}\frac{\partial\theta}{\partial x_{\tau_{i}}}\right],i=1,2,$
$m_{66}\bar{q}_{n}+\mathrm{Kn}\left(m_{67}\frac{\partial\bar{\sigma}_{\langle nn}}{\partial x_{n\rangle}}+m_{68}\frac{\partial\bar{\sigma}_{nj}}{\partial x_{j}}+m_{69}\frac{\partial\theta}{\partial x_{n}}\right)$
(19) $\quad=-\tilde{\chi}\left[-m_{61}(\theta-\theta^{W})+m_{62}\bar{\sigma}_{nn}+m_{63}\mathrm{Kn}\frac{\partial\bar{q}_{j}}{\partial x_{j}}+m_{64}\mathrm{Kn}\frac{\partial\bar{q}_{\langle n}}{\partial x_{n\rangle}}+m_{65}\mathrm{Kn}\frac{\partial v_{\langle n}}{\partial x_{n\rangle}}\right],$
@@
In these equations, $m_{jk}$ are constants depending on the molecular interactions. As examples, we will consider inverse-power-law models in which the repulsive force between two gas molecules is proportional to the $\eta$th power of the distance between them. For such models, the values of the constants $k_{i}$ and $m_{ij}$ are tabulated in the Appendix B for some power indices.